OpenStudy (anonymous):

Factor 3x^2-x-4 and 3x^2-12x+7 Please help...

OpenStudy (reemii):

what about the \(\frac{-b\pm\sqrt\Delta}{2a}\) formula?

OpenStudy (anonymous):

We are using the distributive method....So yeahhh

OpenStudy (reemii):

just use it.

OpenStudy (anonymous):

nd you do that how?

OpenStudy (reemii):

You have seen this formula before? take 3x^2-x-4. \(a=3,b=-1,c=-4\Rightarrow \Delta = (-1)^2 - 4\times 3\times (-4) = 1 + 48 = 49\). the two roots of the first polynomial are: \(\frac{-b\pm\sqrt{49}}{2\times 3}\) .

OpenStudy (anonymous):

Omg I am so confused:/

OpenStudy (reemii):

a=3,b=−1,c=−4 is this part ok?

OpenStudy (anonymous):

Yes

OpenStudy (reemii):

the general form of the polynomial of degree2 is ax^2+bx+c. Yours is 3x^2-x-4. just read a=3,b=−1,c=−4 Then the formula is \(\Delta = b^2-4ac\).

OpenStudy (reemii):

Then put it in the fraction and you find two solutions because of the \(\pm\) . this works if \(\Delta\ge 0\). if \(\Delta<0\) there's no real solution.

OpenStudy (anonymous):

The answers are: (3x − 1)(x + 4) (3x + 4)(x − 1) (3x − 2)(x + 2) (3x − 4)(x + 1) so it is one of those ^^^

OpenStudy (reemii):

what do you find for the roots of the polynomial?

OpenStudy (anonymous):

I am so confused you just have to factor it

OpenStudy (anonymous):

I found the answer now I need help with 4x^2+25x+6

OpenStudy (reemii):

ok. two options, - take all of the possibilities they offer, distribute and you will see which product gives your polynomial. - use the forumla to find the roots (which are (1+7)/6=4/3 and (1-7)/6=-1) therefore the factorization is \(a(x-x_1)(x-x_2) = 3(x-4/3)(x+1)\)

OpenStudy (reemii):

do the same.

OpenStudy (anonymous):

Still confused

OpenStudy (reemii):

how did you do before?

OpenStudy (anonymous):

x=-0.25,x=-6

OpenStudy (anonymous):

Ok multiply a and c if there is no GCF then find numbers that multiply to get what a times c equals but when added equals B